OpenStudy (anonymous):

what am i doing wrong here? Suppose z=x^2siny, x=3s^2+4t^2, y=10st. A. Use the chain rule to find ∂z∂s and ∂z∂t as functions of x, y, s and t.

OpenStudy (anonymous):

i have dx/ds = 6s, dx/dt=8t, dy/ds=10t dy/dt=10s, dz/dx=2xsiny, dz/dy=x^2cosy and an overall dz/ds=dz/dx * dy/ds+dz/dy*dy/ds therefore dz/ds=2xsiny*6x+x^2cosy*10t....webwork says it's incorrect. probably something small and stupid. can someone help?

OpenStudy (anonymous):

@ganeshie8, you've helped me before. can you help me now?

OpenStudy (anonymous):

nevermind. i mistyped my answer. i was right. thanks!

OpenStudy (anonymous):

smart

ganeshie8 (ganeshie8):

chainrule is tricky with more than one variable, but i see you're applying it correctly